how do you do the problem 3x-5=7 and 3x/2=3


Found 2 solutions by solver91311, Cromlix:
Answer by solver91311(24713)



This is two questions, not one; hence you have disregarded the rules for posting and ignored the rules of grammar all in one incorrectly punctuated and incorrectly capitalized sentence. Quite a feat, I must say.

What is it that you do not understand about what you are being asked to do?

What have you tried so far?

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by Cromlix(4381)

I understand that these are two separate problems.
3x - 5 = 7
3x = 7 + 5
3x = 12
x = 12/3
x = 4
3x/2 = 3
Multiply both sides by 2
3x = 6
x = 6/3
x = 2
To check you can always put the answers
back into the original equations:-
3x - 5 = 7
3*4 - 5 = 7
12 - 5 = 7 which is correct
3x/2 = 3
3*2/2 = 3
6/2 = 3 which is correct.
Hope this helps
:-)